Write the equation of the circle that has a diameter whose endpoints are (3,-5) and (-7,3). Write the equation in General Form which is: x2+y2+Dx+Ey+F=0
The "2's" are exponents. I am stumped on this last bit of work I need to know. Can you please help me?
Thanks!

Answer by scott8148(3751)

an equation for a circle centered at (h,k) with a radius of r is (x-h)^2+(y-k)^2=r^2

the center of this circle is the midpoint of the diameter ... ((3-7)/2,(-5+3)/2) or (-2,-1)

the radius is the distance from the center to one of the ends ... r^2=(-2-3)^2+(-1-(-5))^2 ... r^2=25+16=41

so (x+2)^2+(y+1)^2=41 ... x^2+4x+4+y^2+2y+1=41 ... x^2+y^2+4x+2y-36=0